OpenStudy (anonymous):

Hello I'm new to this site and I really need help!! My question is, A circle has a radius of 11 inches and a central angle AOB that measures 45 degrees. What is the area of sector AOB? Use 3.14 for pi and round your answer to the nearest tenth. I really appreciate it if someone could help me!

OpenStudy (anonymous):

A) 47.5 in^2 B) 11.9 in^2 C) 8.6 in ^2 D) 4.3^2

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

I really do not know how to work this site and I can't really see your response

OpenStudy (anonymous):

we can make a proportion, compare angle to angle and circumference to circumference

OpenStudy (anonymous):

Lol and how do you do that I'm like completely lost here?!

OpenStudy (anonymous):

Answer letter B

OpenStudy (anonymous):

$$ \Large \frac{\theta^{\circ}}{360^{\circ}} = \frac{x}{\pi r^2 } $$

OpenStudy (anonymous):

we want to compare angle to angle and area to area

OpenStudy (anonymous):

The directions say to use 3.14 for \( \large \pi\), and radius is 11 $$ \Large \frac{45^{\circ}}{360^{\circ}} = \frac{x}{3.14 \cdot 11^2 } $$

OpenStudy (anonymous):

Thank you both so much!

OpenStudy (anonymous):

now solve for x, cross multiply

OpenStudy (anonymous):

Ok I'm taking a math quiz right now and I don't get Math so can anyone help me with this quiz! Please!